Write the mathematical expression for the reaction quotient, Qc, for each of the following reactions:

(a) CH4(g) + Cl2(g) ⇌ CH3Cl(g) + HCl(g)

(b) 2SO2(g) + O2(g) ⇌ 2SO3(g)

(c) BaSO3(s) ⇌ BaO(s) + SO2(g)

 (d) CH4(g) + 2O2(g) ⇌ CO2(g) + 2H2O(l)

(e) CO2(g) + H2(g) ⇌ CO(g) + H2O(g)



1
Expert's answer
2021-11-13T01:38:08-0500

Qc describes a reaction that is not at the equilibrium.

a)."Qc=\dfrac{[CH_3Cl][HCl]]}{[CH_4][Cl]}"



b.)"Qc=\dfrac{[SO_3]^2}{[SO_2]^2[O_2]}"


c."Qc=\dfrac{[BaO][SO_2]}{[BaSO_3]}"